# Hilbert’s fourth problem in the constant curvature setting

### v1

@@ Abstract.
B. Li is supported by Natural Science Foundation of Zhejiang Province (No. LY23A010012) and Natural Science Foundation of Ningbo (No. 2024J017). W. Zhao is supported by Natural Science Foundation of China (No. 12471045) and Science and Technology Project of Xinjiang Production and Construction Corps (No. 2023CB008-12).
@@ 1.1. Background and motivation
Over the past century, numerous significant results concerning both symmetric and asymmetric distances have been established, as detailed in the monographs by Papadopoulos–Troyanov [25], Pogorelov [26], and others. Nevertheless, the classification problem remains as remarkably challenging. This difficulty stems from the fact that Hilbert’s fourth problem imposes no specific constraints on the distance functions. Among various distances, those induced by families of pseudo-norm functions have attracted considerable attention. Specifically, consider a continuous function $(x,y)\longmapsto\|y\|_{x}$ on the tangent bundle $TM$ of a manifold $M$ that satisfying the following properties for each $x\in M$ and $y,v\in T_{x}M$:
@@
on $\Omega=\{x:1+\lambda|x|^{2}>0\}$. The Finslerian setting, freed from the constraint of quadratic metrics, exhibits a far richer structure. The regularity hypotheses (iv) and (v) permit the extension of sectional curvature to Finsler geometry via the notion of flag curvature. In contrast to the Riemannian case, projective flatness and constant flag curvature are independent conditions in Finsler geometry; neither implies the other. Moreover, even when both conditions are simultaneously satisfied, they admit a vast and largely unexplored family of metrics.
@@
The case of positive constant flag curvature remained open until Bryant’s breakthrough in 1996 [6], who constructed the first non-Riemannian, irreversible, projectively flat Finsler metric with $\mathbf{K}=1$ on $\mathbb{S}^{2}$. Using an algebraic approach, Shen [31] later extended Bryant’s metric to higher-dimensional spheres $\mathbb{S}^{n}$ and observed that such metrics arise from projectively flat metrics on $\mathbb{R}^{n}$.
@@ 1.2. Main results
For the positive curvature case, since $\Omega$ is noncompact, the Bonnet–Myers theorem (cf. Bao–Chern–Shen [1, Theorem 7.7.1]) implies that every projectively flat Finsler manifold $(\Omega,F)$ with $\mathbf{K}=1$ must be incomplete. This motivates the study of its completion. Let $\mathbb{S}^{n}_{+}$ denote the upper hemisphere and $\mathfrak{p}:\mathbb{R}^{n}\rightarrow\mathbb{S}^{n}_{+}\subset\mathbb{R}^{n+1}$ the natural diffeomorphism
@@
It is well-established that $F_{1}$ defined in (1.6) is the unique projectively flat Riemannian metric on $\mathbb{R}^{n}$ with $\mathbf{K}=1$ and the pull-back metric $(\mathfrak{p}^{-1})^{*}F_{1}$ coincides precisely with the standard sphere metric. Thus, Theorem 1.5 is natural in the Riemannian case. For the non-Riemannian metrics, however, the assumption that $(\mathfrak{p}^{-1})^{*}F$ can be extended to $\partial\mathbb{S}^{n}_{+}$ is necessary. Indeed, there exist projectively flat Finsler metrics on $\mathbb{R}^{n}$ with $\mathbf{K}=1$ for which this extension fails. An example and further discussion are provided in Remark 5.8.
@@ 2.1. Positively homogeneous functions
which implies $\varphi(\mathbf{0})=0$ when $\kappa>0$. If we further assume $\varphi|_{\operatorname{\mathbb{R}_{\,\circ}^{\it n}}}\in C^{1}(\operatorname{\mathbb{R}_{\,\circ}^{\it n}})$, then Euler’s theorem (cf. Bao–Chern–Shen [1, Theorem 1.2.1]) yields
@@
The following definition is due to Busemann–Mayer [9, p. 181]:
@@
A quasi-regular function is always continuous due to its convexity. Moreover, every compact convex set in $\mathbb{R}^{n}$ corresponds to a quasi-regular function (cf. Hörmander [18, Theorem 2.1.23]). In addition, the following result holds (cf. [9, pp. 178–179]).
@@
where $\varphi$ is a positively $1$-homogeneous function on $\mathbb{R}^{n}$. Equation (2.3) is closely related to projectively flat Finsler metrics (cf. [23, 31]). For instance, the celebrated Funk metric arises as a solution, while the Hilbert metric is obtained by symmetrizing the former.
@@
In particular, Shen proved the local uniqueness of solutions to (2.3) (cf. [31, Lemma 5.1]).
@@
$0\geq\varphi(-x)=\frac{\Phi(x,-x)}{1-\Phi(x,-x)}>-1,\quad\forall x\in\mathcal{D}(\Phi).$
The proof is completed. ∎
@@ 2.2. Elements from Finsler geometry
This subsection recalls essential definitions and properties from Finsler geometry; for comprehensive treatments, we refer to Bao–Chern–Shen [1], Ohta [24] and Shen [32, 30].
@@
Every pseudo-norm is quasi-regular and every Minkowski norm is regular. Furthermore, the class of weak Minkowski norms is strictly weaker, as there exist examples that are not Minkowski norms (see [1, Exercise 1.2.7]). A fundamental example of Minkowski norm on $\mathbb{R}^{n}$ is the Randers norm $\psi(y):=|y|+\langle a,y\rangle$, where $a$ is a fixed vector with $|a|<1$.
@@
where
$R^{i}_{\ k}=2G^{i}_{x^{k}}-y^{j}G^{i}_{x^{j}y^{k}}+2G^{j}G^{i}_{y^{j}y^{k}}-G^{i}_{y^{j}}G_{y^{k}}^{j}.$
A Finsler metric $F$ is of scalar flag curvature if the curvature depends only on the position and direction, namely, $\mathbf{K}(y;\Pi)=\mathbf{K}(x,y)$. It is of constant flag curvature $k$ if $\mathbf{K}\equiv k$. For a Riemannian metric, the flag curvature specializes to the sectional curvature.
@@
Let $(r,y)$ be the polar coordinate system around some point $o\in M$. That is, if $x=(r,y)$, then $x=\exp_{o}(ry)$ with $r=r(x)=d_{F}(o,x)$ and $y=y(x)\in S_{o}M$ (cf. Zhao and Shen [34, Section 3]). Hence, for any $y\in S_{o}M$ and $r\in(0,i_{y})$,
@@
Moreover, it follows from [30, Lemma 3.2.3] that $F(\nabla r)=F^{*}({\operatorname{d}}r)=1$ for $\operatorname{\mathfrak{m}}$-a.e. $x\in M$.
@@
The following volume comparison result is essential to our analysis of Sobolev spaces in Section 6 (cf. [34, Theorem 3.6]).
@@
We conclude this subsection by recalling the integral of distortion
@@ 2.3. Projectively flat Finsler metrics
In this subsection, we present the basic properties and some new results of projectively flat Finsler metrics, which are the regular pseudo-norm functions concerning Hilbert’s fourth problem. See Berwald [4, 5], Funk [14], Papadopoulos–Troyanov [25], Shen [32], etc. for further details.
@@
Additionally, we have the following completeness result (cf. Kristály–Li–Zhao [20, Proposition 6.1]).
@@
It is natural to ask whether $\varphi$ is a Minkowski norm in Lemma 2.21. Unfortunately, this may fail; see [1, Exercise 1.2.7] for an example. In fact, $\varphi$ is a Minkowski norm if and only if $\partial\Omega$ is strongly convex, i.e., its Gaussian curvature is positive everywhere (cf. [25, p. 25]).
@@ 3.1. Global characterization of manifolds
This subsection begins by recalling a local characterization theorem due to Shen [31, Theorem 1.3] and Li [23, Theorem 1.1] for the sufficiency and necessity, respectively.
@@ 3.2. Global existence of metrics
Moreover, if the matrix $([F^{2}]_{y^{i}y^{j}}(x,y))$ is positive definite, then by [1, (1.2.9)] we have
@@
This example motivates a fundamental question: under what conditions does the function $F$ associated with a pair $(\psi,\phi)$ fail to be a Finsler metric on the entire $\Omega$? To answer this, we prove the following key lemma in Appendix A.
@@
The next lemma follows directly from Rademacher’s Theorem (cf. Evans [12, Theorem 6]) and Grisvard [15, Corollary 1.2.2.3].
@@ 4.1. Global characterization of manifolds
The starting point for our global analysis is the following local structure theorem. The sufficiency and necessity were established by Shen [31, Theorem 1.2] and Li [23, Theorem 1.2], respectively, while the reversible case was treated in Cheng–Li [11, Theorem 1.1, Theorem 3.4].
@@
By the Bonnet–Myers theorem (cf. [1, Theorem 7.7.1]), every projectively flat Finsler manifold $(\Omega,F)$ of constant positive flag curvature cannot be forward complete because $\Omega$ is noncompact. Therefore, it suffices to consider the case of non-positive constant flag curvature, which follows directly from Corollaries 3.5 and 4.9. ∎
@@ 4.2. Global existence of metrics
By Theorem 2.16, the answer is affirmative when $\phi=\psi$, in which case $F=\Theta/2$ for the unique Funk metric $\Theta$ on $\Omega$. A Hilbert metric also exists globally on $\Omega$ as the symmetrization of $\Theta$. For other pairs $(\psi,\phi)$, however, phenomena analogous to the $\mathbf{K}=0$ case occur. We now introduce the notation used to study this question.
@@
The forward implication is immediate from the definitions. The converse follows directly from Corollary 4.11. ∎
@@
which forces $\theta=\mathbf{0}$. This contradicts the assumption that $\theta\neq\mathbf{0}$. Therefore, no such point $\breve{x}$ exists, and $U=\Omega=\operatorname{\mathscr{O}}$.
@@ 5.1. Distance and diameter
By the Bonnet–Myers theorem, any forward complete Finsler manifold with $\mathbf{K}=1$ is compact and has diameter $\leq\pi$. It follows that a projectively flat Finsler manifold $(\mathbb{R}^{n},F)$ with $\mathbf{K}=1$ cannot be forward complete. Nevertheless, Theorem 5.3 shows that such a manifold still attains the maximum possible diameter $\pi$. In the next subsection, we prove that under a natural necessary condition, the completion of such a manifold is a sphere.
@@ 5.2. Completion of manifolds
The maximum diameter theorem in the Riemannian setting states that a complete Riemannian manifold with sectional curvature $1$ and diameter $\pi$ must be a sphere. In view of Theorem 5.3, a natural question arises: is the completion of a projectively flat Finsler manifold $(\mathbb{R}^{n},F)$ with $\mathbf{K}=1$ also a sphere? The answer is affirmative under a natural necessary condition.
@@
For any $(w,V)\in S\mathbb{S}^{n}$ with $V^{n+1}\neq 0$, the injectivity radius satisfies $i_{V}=i_{(w,V)}=\pi$. For the remaining case where $(w,V)\in S\mathbb{S}^{n}$ with $w\in\partial\mathbb{S}^{n}_{+}$ and $V^{n+1}=0$, choose a sequence $(V_{\alpha})_{\alpha}\subset S_{w}\mathbb{S}^{n}$ converging to $V$. The continuity of the injectivity radius (cf. [1, Proposition 8.4.1]) then yields $i_{V}=\lim_{\alpha\rightarrow+\infty}i_{V_{\alpha}}=\pi$. Hence, the injectivity radius is $\pi$ everywhere. Consequently, $d_{\mathscr{F}}(w,-w)=\pi$ for all $w\in\mathbb{S}^{n}$, since there always exists a geodesic (a great semicircle) not entirely contained in $\partial\mathbb{S}^{n}_{+}$ from $w$ to w-w of length $\pi$.
@@
Conclusion (i) follows immediately from the Bonnet–Myers theorem (cf. [1, Theorem 7.7.1]) combined with the reverse metric, while Conclusions (ii) and (iii) are direct consequences from Theorems 5.3 and 5.6. ∎
We conclude this subsection with concrete examples illustrating Theorem 5.6. Using (LABEL:K=1spherical_x), a direct computation yields
@@
Using (LABEL:K=1spherical_x), (5.24), and (5.25), the pull-back metric on $\mathbb{S}^{n}_{+}$ is
@@
Example 5 (Bryant’s metric[7, 31]).

### v2

@@ 1.1. Background and motivation
Over the past century, numerous significant results concerning both symmetric and asymmetric distances have been established, as detailed in the works of Busemann [16], Pogorelov [40], Szabó [49], Álvarez Paiva [1], Papadopoulos–Troyanov [39], and others. Nevertheless, the classification problem remains as remarkably challenging. This difficulty stems from the fact that Hilbert’s fourth problem imposes no specific constraints on the distance functions. Among various distances, those induced by families of pseudo-norm functions have attracted considerable attention. Specifically, consider a continuous function $(x,y)\longmapsto\|y\|_{x}$ on the tangent bundle $TM$ of a manifold $M$ that satisfying the following properties for each $x\in M$ and $y,v\in T_{x}M$:
@@
on $\Omega=\{x:1+\lambda|x|^{2}>0\}$. The Finslerian setting, freed from the constraint of quadratic metrics, exhibits a far richer structure. The regularity hypotheses (iv) and (v) permit the extension of sectional curvature to Finsler geometry via the notion of flag curvature. In contrast to the Riemannian case, projective flatness and constant flag curvature are independent conditions in Finsler geometry; neither implies the other. The classification of Finsler metrics of constant flag curvature is far from understood. While a complete classification has been achieved for Randers metrics (cf. Bao–Robles–Shen [5]), comprehensive characterizations for other classes have been more limited. Moreover, even when both conditions are simultaneously satisfied, they admit a vast and largely unexplored family of metrics.
@@
The case of positive constant flag curvature remained open until Bryant’s breakthrough in 1996–1997 [9, 10], who constructed the first non-Riemannian, irreversible, projectively flat Finsler metric with $\mathbf{K}=1$ on $\mathbb{S}^{2}$. He later discussed broader generalizations in [11]. Building on this foundation, Shen [48] used an algebraic approach to extend Bryant’s metric to higher-dimensional spheres $\mathbb{S}^{n}$, observing that such metrics arise from projectively flat metrics on $\mathbb{R}^{n}$.
@@ 1.2. Main results
For the positive curvature case, since $\Omega$ is noncompact, the Bonnet–Myers theorem (cf. Bao–Chern–Shen [4, Theorem 7.7.1]) implies that every projectively flat Finsler manifold $(\Omega,F)$ with $\mathbf{K}=1$ must be incomplete. This motivates the study of its completion. Let $\mathbb{S}^{n}_{+}$ denote the upper hemisphere and $\mathfrak{p}:\mathbb{R}^{n}\rightarrow\mathbb{S}^{n}_{+}\subset\mathbb{R}^{n+1}$ the natural diffeomorphism
@@
It is well-known that $F_{1}$ in (1.6) is the unique projectively flat Riemannian metric on $\mathbb{R}^{n}$ with $\mathbf{K}=1$, and its pull-back $(\mathfrak{p}^{-1})^{*}F_{1}$ is the standard sphere metric. Thus, Theorem 1.5 is natural for Riemannian metrics. For non-Riemannian ones, however, the assumption that $(\mathfrak{p}^{-1})^{*}F$ extends to $\partial\mathbb{S}^{n}_{+}$ is essential, as there exist projectively flat Finsler metrics with $\mathbf{K}=1$ for which this fails; see Remark 5.8.
Theorem 1.5/(iii) implies that all geodesics are closed, of length $2\pi$, and mutually intersecting. This configuration falls into the first case of Bryant–Foulon–Ivanov–Matveev–Ziller [13, Theorem 2], but does not belong to the class of Finsler metrics constructed in Katok [33]. For further related studies on closed geodesics in Finsler geometry, we refer to Bangert–Long [2], Duan–Long–Wang [20], and Duan–Qi [21], and references therein.
@@
Acknowledgements. The authors are grateful to Robert L. Bryant for his guidance regarding relevant references. The first author is supported by Natural Science Foundation of Zhejiang Province (No. LY23A010012) and Natural Science Foundation of Ningbo (No. 2024J017). The second author is supported by Natural Science Foundation of China (No. 12471045) and Science and Technology Project of Xinjiang Production and Construction Corps (No. 2023CB008-12).
@@ 2.1. Positively homogeneous functions
which implies $\varphi(\mathbf{0})=0$ when $\kappa>0$. If we further assume $\varphi|_{\operatorname{\mathbb{R}_{\,\circ}^{\it n}}}\in C^{1}(\operatorname{\mathbb{R}_{\,\circ}^{\it n}})$, then Euler’s theorem (cf. Bao–Chern–Shen [4, Theorem 1.2.1]) yields
@@
The following definition is due to Busemann–Mayer [17, p. 181]:
@@
A quasi-regular function is always continuous due to its convexity. Moreover, every compact convex set in $\mathbb{R}^{n}$ corresponds to a quasi-regular function (cf. Hörmander [30, Theorem 2.1.23]). In addition, the following result holds (cf. [17, pp. 178–179]).
@@
where $\varphi$ is a positively $1$-homogeneous function on $\mathbb{R}^{n}$. This equation is closely related to projectively flat Finsler metrics (cf. [37, 48]). For instance, the celebrated Funk metric arises as a solution, while the Hilbert metric is obtained by symmetrizing the former. Equation (2.3) can also be regarded as a special case of Zermelo deformation when $\varphi$ is a Minkowski norm, a framework known to have profound connections with flag curvature in Finsler geometry, as explored by Foulon [24], Huang–Mo [32], Bryant–Huang–Mo [12], and Foulon–Matveev [25]. In this paper, however, we focus on the more general case where $\varphi$ is merely positively $1$-homogeneous or quasi-convex, which naturally involves the underlying manifold’s boundary.
@@
In particular, Shen proved the local uniqueness of solutions to (2.3) (cf. [48, Lemma 5.1]).
@@
$0\geq\varphi(-x)=\frac{\Phi(x,-x)}{1-\Phi(x,-x)}>-1,\quad\forall x\in\mathcal{D}(\Phi).\qed$
@@ 2.2. Elements from Finsler geometry
This subsection recalls essential definitions and properties from Finsler geometry; for details, we refer to Bao–Chern–Shen [4], Ohta [38] and Shen [47, 46].
@@
Every pseudo-norm is quasi-regular and every Minkowski norm is regular. Furthermore, the class of weak Minkowski norms is strictly weaker, as there exist examples that are not Minkowski norms (see [4, Exercise 1.2.7]). A fundamental example of Minkowski norm on $\mathbb{R}^{n}$ is the Randers norm $\psi(y):=|y|+\langle a,y\rangle$, where $a$ is a fixed vector with $|a|<1$.
@@
where $R^{i}_{\ k}:=2G^{i}_{x^{k}}-y^{j}G^{i}_{x^{j}y^{k}}+2G^{j}G^{i}_{y^{j}y^{k}}-G^{i}_{y^{j}}G_{y^{k}}^{j}.$ A Finsler metric $F$ is of scalar flag curvature if the curvature depends only on the position and direction, namely, $\mathbf{K}(y;\Pi)=\mathbf{K}(x,y)$. It is of constant flag curvature $k$ if $\mathbf{K}\equiv k$. For a Riemannian metric, the flag curvature specializes to the sectional curvature.
@@
Let $(r,y)$ be the polar coordinate system around some point $o\in M$. That is, if $x=(r,y)$, then $x=\exp_{o}(ry)$ with $r=r(x)=d_{F}(o,x)$ and $y=y(x)\in S_{o}M$ (cf. Zhao and Shen [52, Section 3]). Hence, for any $y\in S_{o}M$ and $r\in(0,i_{y})$,
@@
Moreover, it follows from [46, Lemma 3.2.3] that $F(\nabla r)=F^{*}({\operatorname{d}}r)=1$ for $\operatorname{\mathfrak{m}}$-a.e. $x\in M$.
@@
Moreover, $\nu_{o}(S_{o}M)$ plays an important role in the Gauss-Bonnet formula; see Bao–Chern [3] and Shen [44].
The following volume comparison result is essential to our analysis of Sobolev spaces in Section 6 (cf. [52, Theorem 3.6]).
@@
Further comparison results can be found in the works of Shen [45], Burago–Ivanov [15], and Wu–Xin [51], and references therein. We conclude this subsection by recalling the integral of distortion
@@ 2.3. Projectively flat Finsler metrics
This subsection covers basic properties and new results on projectively flat Finsler metrics, which are the regular pseudo-norm functions concerning Hilbert’s fourth problem. See Berwald [7, 8], Funk [26], Papadopoulos–Troyanov [39], and Shen [47] for further details.
@@
Additionally, we have the following completeness result (cf. Kristály–Li–Zhao [34, Proposition 6.1]).
@@
It is natural to ask whether $\varphi$ is a Minkowski norm in Lemma 2.21. Unfortunately, this may fail; see [4, Exercise 1.2.7] for an example. In fact, $\varphi$ is a Minkowski norm if and only if $\partial\Omega$ is strongly convex, i.e., its Gaussian curvature is positive everywhere (cf. [39, p. 25]).
@@ 3.1. Global characterization of manifolds
This subsection begins by recalling a local characterization theorem due to Shen [48, Theorem 1.3] and Li [37, Theorem 1.1] for the sufficiency and necessity, respectively.
@@ 3.2. Global existence of metrics
Moreover, if the matrix $([F^{2}]_{y^{i}y^{j}}(x,y))$ is positive definite, then by [4, (1.2.9)] we have
@@
This example motivates a fundamental question:
Under what conditions does the function $F$ associated with a pair $(\psi,\phi)$ fail to be a Finsler metric on the entire $\Omega$?
To answer this, we prove the following key lemma in Appendix A.
@@
The next lemma follows directly from Rademacher’s Theorem (cf. Evans [22, Theorem 6]) and Grisvard [27, Corollary 1.2.2.3].
@@ 4.1. Global characterization of manifolds
The starting point for our global analysis is the following local structure theorem. The sufficiency and necessity are due to Shen [48, Theorem 1.2] and Li [37, Theorem 1.2], respectively, while the reversible case was treated in Cheng–Li [18, Theorems 1.1 and 3.4].
@@
By the Bonnet–Myers theorem (cf. [4, Theorem 7.7.1]), every projectively flat Finsler manifold $(\Omega,F)$ of constant positive flag curvature cannot be forward complete because $\Omega$ is noncompact. Therefore, it suffices to consider the case of non-positive constant flag curvature, which follows directly from Corollaries 3.5 and 4.9. ∎
@@ 4.2. Global existence of metrics
By Theorem 2.16, the answer is affirmative when $\phi=\psi$, in which case $F=\Theta/2$ for the unique Funk metric $\Theta$ on $\Omega$, with its symmetrization yielding the Hilbert metric. For other pairs $(\psi,\phi)$, however, phenomena analogous to the $\mathbf{K}=0$ case occur. We now introduce the notation for this question.
@@
The necessity is immediate from the definitions, and the sufficiency follows from Corollary 4.11. ∎
@@
which forces $\theta=\mathbf{0}$, contradicting the assumption that $\theta\neq\mathbf{0}$. Thus, no such point $\breve{x}$ exists, so $U=\Omega=\operatorname{\mathscr{O}}$.
@@ 5.1. Distance and diameter
By the Bonnet–Myers theorem, any forward complete Finsler manifold with $\mathbf{K}=1$ is compact and has diameter $\leq\pi$. Hence, a projectively flat Finsler manifold $(\mathbb{R}^{n},F)$ with $\mathbf{K}=1$ cannot be forward complete. Theorem 5.3 yields the maximal diameter $\pi$, and we next show its completion is a sphere under a natural condition.
@@ 5.2. Completion of manifolds
The maximum diameter theorem in the Riemannian setting states that a complete Riemannian manifold with sectional curvature $1$ and diameter $\pi$ must be a sphere. In view of Theorem 5.3, a natural question arises:
Is the completion of a projectively flat Finsler manifold $(\mathbb{R}^{n},F)$ with $\mathbf{K}=1$ also a sphere?
The answer is affirmative under a natural necessary condition.
@@
For any $(w,V)\in S\mathbb{S}^{n}$ with $V^{n+1}\neq 0$, the injectivity radius satisfies $i_{V}=i_{(w,V)}=\pi$. For the remaining case where $(w,V)\in S\mathbb{S}^{n}$ with $w\in\partial\mathbb{S}^{n}_{+}$ and $V^{n+1}=0$, choose a sequence $(V_{\alpha})_{\alpha}\subset S_{w}\mathbb{S}^{n}$ converging to $V$. The continuity of the injectivity radius (cf. [4, Proposition 8.4.1]) then yields $i_{V}=\lim_{\alpha\rightarrow+\infty}i_{V_{\alpha}}=\pi$. Hence, the injectivity radius is $\pi$ everywhere. Consequently, $d_{\mathscr{F}}(w,-w)=\pi$ for all $w\in\mathbb{S}^{n}$, since there always exists a geodesic (a great semicircle) not entirely contained in $\partial\mathbb{S}^{n}_{+}$ from $w$ to w-w of length $\pi$.
@@
Conclusion (i) follows immediately from the Bonnet–Myers theorem (cf. [4, Theorem 7.7.1]) combined with the reverse metric, while Conclusions (ii) and (iii) are direct consequences from Theorems 5.3 and 5.6. ∎
We conclude this subsection with concrete examples illustrating Theorem 5.6. Using (5.17), a direct computation yields
@@
Using (5.17), (5.24), and (5.25), the pull-back metric on $\mathbb{S}^{n}_{+}$ is
@@
Example 5 (Bryant’s metric[9, 10, 48]).